### v1

# Block-Diagonal Orthogonal Relation and Matrix Entity for Knowledge Graph Embedding

Abstract
The primary aim of Knowledge Graph embeddings (KGE) is to learn low-dimensional representations of entities and relations for predicting missing facts. While rotation-based methods like RotatE Sun et al. (2019) and QuatE Zhang et al. (2019) perform well in KGE, they face two challenges: limited model flexibility requiring proportional increases in relation size with entity dimension, and difficulties in generalizing the model for higher-dimensional rotations. To address these issues, we introduce OrthogonalE, a novel KGE model employing matrices for entities and block-diagonal orthogonal matrices with Riemannian optimization for relations. This approach enhances the generality and flexibility of KGE models. The experimental results indicate that our new KGE model, OrthogonalE, is both general and flexible, significantly outperforming state-of-the-art KGE models while substantially reducing the number of relation parameters. Our code is available at OrthogonalE
[…]
Block-Diagonal Orthogonal Relation and Matrix Entity for Knowledge Graph Embedding
Yihua Zhu1,2    Hidetoshi Shimodaira1,2 1Kyoto University   2RIKEN zhu.yihua.22h@st.kyoto-u.ac.jp, shimo@i.kyoto-u.ac.jp
1 Introduction
The fundamental elements of knowledge graphs (KGs) are factual triples, each represented as $(h,r,t)$, indicating a relationship $r$ between head entity $h$ and tail entity $t$. Notable examples include Freebase Bollacker et al. (2008), Yago Suchanek et al. (2007), and WordNet Miller (1995). KGs have practical applications in diverse fields such as question-answering Hao et al. (2017), information retrieval Xiong et al. (2017), recommender systems Zhang et al. (2016), and natural language processing Yang and Mitchell (2019), garnering considerable interest in academic and commercial research.
Addressing the inherent incompleteness of KGs, link prediction has become a pivotal area of focus. Recent research Bordes et al. (2013); Trouillon et al. (2016) has extensively leveraged Knowledge Graph Embedding (KGE) techniques, aiming to learn compact, low-dimensional representations of entities and relations. These approaches, marked by scalability and efficiency, have shown proficiency in modeling and deducing KG entities and relations from existing facts.
Recently, rotation-based KGE methods have achieved notable success in the field. For instance, RotatE Sun et al. (2019) conceptualizes relations as 2D rotations, while QuatE Zhang et al. (2019) employs 3D rotations. Essentially, as illustrated in Fig. 1 , both operate by multiplying the relation matrix $\mathbf{e}_{R}^{1}\in\mathbb{R}^{n\times n}$ composed of numerous small rotation blocks $\mathbf{B}_{i}\in\mathbb{R}^{d\times d}$ (RotatE: $\mathbb{R}^{2\times 2}$, QuatE: $\mathbb{R}^{3\times 3}$) with the head entity vector $\mathbf{e}_{h}^{1}\in\mathbb{R}^{n}$.
Figure 1: Fundamental operations and inherent challenges of rotation-based KGE models.
Nevertheless, these approaches encounter two primary issues as shown in Fig. 1. First, the model’s lack of flexibility necessitates increasing the overall relation matrix ($\mathbf{e}_{R}^{1}\in\mathbb{R}^{n\times n}\rightarrow\mathbf{e}_{R}^{2}\in\mathbb{R}^{(n+l)\times(n+l)}$) to meet entity dimension requirements ($\mathbf{e}_{h}^{1}\in\mathbb{R}^{n}\rightarrow\mathbf{e}_{h}^{2}\in\mathbb{R}^{n+l}$) for better represent entities, which often leads to redundancy and inefficiency in representing relations. Second, exploring high-dimensional rotational KGE models is challenging due to the significant computational demands and complexity of rotations in higher dimensions ($\mathbf{B}_{i}:\mathbb{R}^{2\times 2},\mathbb{R}^{3\times 3}\rightarrow\mathbb{R}^{d\times d},d>3$), such as 4D and 5D rotation. This restricts the development of more generalized and higher-dimensional rotation KGE approaches.
To overcome these issues, we propose a highly general and flexible KGE model named OrthogonalE as shown in Fig. 3. Firstly, by transforming entity vectors $\mathbf{e}_{h}\in\mathbb{R}^{n}$ into matrices $\mathbf{e}_{H}\in\mathbb{R}^{n\times m}$, we control the entity dimension through variable $m$, avoiding unnecessary expansion of the relation size. Secondly, leveraging the concept that rotation matrices are orthogonal, we replace rotation matrices $\mathbf{B}_{i}$ with orthogonal matrices $\mathbf{X}_{i}\in\mathbb{R}^{d\times d}$ of adaptable dimensions $d$, facilitating the exploration of higher-dimensional block-diagonal orthogonal matrix models. Lastly, for effective optimization, we employ Riemannian optimization for the relation matrix $\mathbf{e}_{R}\in\mathbb{R}^{n\times n}$ and Stochastic Gradient Descent (SGD) for the entity matrix $\mathbf{e}_{V}\in\mathbb{R}^{n\times m}$.
We evaluate the new model on two KGE datasets including WN18RR Dettmers et al. (2018), FB15K-237 Toutanova and Chen (2015). The experimental results demonstrate that our new KGE model, OrthogonalE, is both general and flexible, significantly outperforming state-of-the-art KGE models while substantially reducing the number of relation parameters.
2 Related Work
Knowledge Graph Embedding
Translation-based approaches are prominent in KGE, notably TransE Bordes et al. (2013), which interprets relations as vector translations. TransH Wang et al. (2014), TransR Lin et al. (2015), and TransD Ji et al. (2015) represent extensions of the translation-based method, building upon the foundational approach of TransE. ComplExTrouillon et al. (2016) advances this by embedding entities and relations in a complex space and using the Hermitian product for modeling antisymmetric patterns, a technique that has proven highly effective in learning KG representations. Inspired by ComplEx, RotatE Sun et al. (2019) then innovated by treating relations as rotations in a complex vector space, capable of capturing varied relation patterns like symmetry and inversion. Following this, QuatE Zhang et al. (2019) employed quaternion operations (3D rotations) for even better expressiveness than RotatE.
However, considering the two major disadvantages of rotation-based methods mentioned in the Introduction 1, there is a need to refine our model to make it more general and flexible.
Optimization on the orthogonal manifold
In optimization on the orthogonal manifold, transitioning from $X^{t}$ to $X^{t+1}$ while remaining on the manifold necessitates a method known as retraction Absil and Malick (2012). Prior research has effectively adapted several standard Euclidean function minimization algorithms to Riemannian manifolds. Notable examples include gradient descent (Absil et al. (2008); Zhang and Sra (2016)), second-order quasi-Newton methods (Absil et al. (2007); Qi et al. (2010)), and stochastic approaches Bonnabel (2013), crucial in deep neural network training.
Riemannian Optimization has concurrently advanced in deep learning domains, particularly in CNNs and RNNs. Cho and Lee (2017) innovatively substituted CNN’s Batch Normalization layers with Riemannian optimization on the Grassmann manifold for parameter normalization. Additionally, significant strides in stabilizing RNN training have been made by Vorontsov et al. (2017), Wisdom et al. (2016), Lezcano-Casado and Martınez-Rubio (2019), and Helfrich et al. (2018), through the application of Riemannian optimization to unitary matrices.
As this paper primarily focuses on KGE, we do not delve deeply into Riemannian optimization. Instead, we utilize the retraction with exponential map method for iterative optimization, sourced from the Geoopt Kochurov et al. (2020).
3 Problem Formulation and Background
We present the KGE problem and describe Optimization on the orthogonal manifold before our approach part.
3.1 Knowledge Graph Embedding
In a KG consisting of fact triples $(h,r,t)\in\mathcal{E}\subseteq\mathcal{V}\times\mathcal{R}\times\mathcal{V}$, with $\mathcal{V}$ and $\mathcal{R}$ denoting entity and relation sets, the objective of KGE is to map entities $v\in\mathcal{V}$ to $k_{\mathcal{V}}$-dimensional embeddings $\mathbf{e}_{v}$, and relations $r\in\mathcal{R}$ to $k_{\mathcal{R}}$-dimensional embeddings $\mathbf{e}_{r}$.
A scoring function $s:\mathcal{V}\times\mathcal{R}\times\mathcal{V}\rightarrow\mathbb{R}$ evaluates the difference between transformed and target entities, quantified as a Euclidean distance:
$d^{E}\left(\mathbf{x},\mathbf{y}\right)=\left\|\mathbf{x}-\mathbf{y}\right\|$
3.2 Optimization on the orthogonal manifold
Figure 2: Abstract representation of Riemannian gradient descent iteration on orthogonal manifold
In optimization on the orthogonal manifold, the core problem is formulated as:
$\min_{X\in\mathcal{O}_{d}}f(X),$ (1)
Here, $f$ is a differentiable function mapping elements of $\mathbb{R}^{d\times d}$ to $\mathbb{R}$, and the orthogonal manifold $\mathcal{O}_{d}$ is defined as $\mathcal{O}_{d}\triangleq\left\{X\in\mathbb{R}^{d\times d}\mid XX^{\top}=I_{d}\right\}$. Moreover, the tangent space at $X$, denoted by $\mathcal{T}_{X}$, is the set $\mathcal{T}_{X}=\left\{\xi\in\mathbb{R}^{d\times d}\mid\xi X^{\top}+X\xi^{\top}=0\right\}$.
To address problem 1 more efficiently, recent studies suggest optimization on the orthogonal manifold with retractions as an effective approach Ablin and Peyré (2022). In this work, we primarily employ the retraction with exponential map method for iterative optimization, as illustrated in Figure 2. The key iteration formula for this method is:
$X^{t+1}=\text{Exp}_{X^{t}}(-\eta\text{Grad}f(X^{t})),$ (2)
Where $t$ indexes the iteration steps, $\text{Exp}_{X^{t}}(\xi)$ denotes the exponential map, and $\eta$ represents the learning rate. $\text{Grad}f(\cdot)$ is the Riemannian gradient. Subsequent sections will delve into the computation of $\text{Exp}_{X^{t}}(\xi)$ and $\text{Grad}f(\cdot)$.
The exponential map allows movement in a specified direction on the manifold. Starting from $X$ with initial velocity $\xi$, the exponential map for the orthogonal matrices manifold is represented by Massart and Abrol (2022):
$\text{Exp}_{X}(\xi)=X\text{expm}(X^{\top}\xi),\forall\xi\in\mathcal{T}_{X},$
Where $\text{expm}()$ denotes the matrix exponential.
On the orthogonal manifold, the Riemannian gradient $\text{Grad}f(\cdot)$ is calculated as Absil et al. (2008):
$\text{Grad}f(X)=P_{\mathcal{T}_{X}}(\bigtriangledown f(X)),$
Where $\bigtriangledown f(X)$ is Euclidean gradient of $f(X)$, and the calculation formula for $P_{\mathcal{T}_{X}}(\cdot)$ is:
$P_{\mathcal{T}_{X}}(Y)=X(\frac{X^{\top}Y-Y^{\top}X}{2}),Y\in\mathbb{R}^{d\times d}$
4 Approach
Figure 3: Diagram of the OrthogonalE approach.
Our approach aims to acquire a flexible and general KGE model with employing matrices for entities and block-diagonal orthogonal matrices with Riemannian optimization for relations.
4.1 Orthogonal Matrices for Relations
RotatE (2D rotation) and QuatE (3D rotation) inherently employ fixed rotation matrices ($\mathbf{B}_{i}\in\mathbb{R}^{d\times d},d\leqslant 3$), as depicted in Figure 1. However, the exploration of high-dimensional rotational KGE models ($\mathbf{B}_{i}\in\mathbb{R}^{d\times d},d>3$) is limited by substantial computational demands and the complexity of rotations in higher dimensions, which hinders the advancement of more generalized, high-dimensional rotational KGE models. To address this, we exploit the orthogonality of rotation matrices, substituting rotation matrices($\mathbf{B}_{i}\in\mathbb{R}^{d\times d}$) with orthogonal matrices ($\mathbf{X}_{i}\in\mathbb{R}^{d\times d}$) of corresponding dimensions $d$. Consequently, our relation embedding ($\mathbf{e}_{R}\in\mathbb{R}^{n\times n}$) are composed of $n/d$ block-diagonal orthogonal matrices $\mathbf{X}_{i}$ as illustrated in Figure 3:
$\mathbf{e}_{R}=diag(\mathbf{X}_{1},\mathbf{X}_{2},...,\mathbf{X}_{n/d})$ (3)
Where the number of relation parameters is $\frac{d(d-1)}{2}*\frac{n}{d}=\frac{(d-1)n}{2}$. And this aspect allows OrthogonalE to gain generality, adapting to datasets with diverse complexities by modifying the block-diagonal matrices’ dimension dd."
4.2 Matrices Representation for Entities
To enhance OrthogonalE’s flexibility, we aim to regulate entity dimension using variable $m$ and transform entity vectors $\mathbf{e}_{v}\in\mathbb{R}^{n}$ into matrices $\mathbf{e}_{V}\in\mathbb{R}^{n\times m}$ as shown in Figure 3, thus preventing unnecessary expansion of the relation size. This part allows OrthogonalE to acquire flexibility, adapting to diverse datasets with varying relation and entity parameters, rather than indiscriminately increase both. And the number of entity parameters is nmn*m.
4.3 Scoring function and Loss
We utilize the Euclidean distance between the transformed head entity $\mathbf{e}_{R}\cdot\mathbf{e}_{H}$ and the tail entity $\mathbf{e}_{T}$ as the scoring function:
$s(h,r,t)=-d^{E}\left(\mathbf{e}_{R}\cdot\mathbf{e}_{H},\mathbf{e}_{T}\right)+\!b_{h}\!+\!b_{t}$ (4)
Here, $b_{v}(v\in\mathcal{V})$ denotes the entity bias, incorporated as a margin in the scoring function, following methodologies from Tifrea et al. (2018); Balazevic et al. (2019). Furthermore, we opt for uniform selection of negative samples for a given triple $(h,r,t)$ by altering the tail entity, rather than employing alternative negative sampling techniques. The loss function defined as follows:
$L=\sum_{t^{\prime}}\log\left(1+\exp\left(y_{t^{\prime}}\cdot s\left(h,r,t^{\prime}\right)\right)\right)$ (5)
[…]
4.4 Optimization
Traditional KGE models train and optimize relations and entities jointly. In contrast, our study aims for more effective optimization of the relation embeddings’ block-diagonal orthogonal matrices $\mathbf{X}_{i}\in\mathbb{R}^{d\times d}$ by separately optimizing relations and entities, utilizing Riemannian optimization for the relation matrix $\mathbf{e}_{R}\in\mathbb{R}^{n\times n}$ and SGD for the entity matrix $\mathbf{e}_{V}\in\mathbb{R}^{n\times m}$.
Initially, when optimizing relations, all entity parameters are fixed, rendering the entity embeddings analogous to the function $f(\cdot)$ in the problem 1. Notably, each orthogonal matrix $\mathbf{X}_{i}$ within the relation embedding $\mathbf{e}_{R}$ optimized by individual Riemannian optimization using RiemannianAdam Kochurov et al. (2020), which is a Riemannian version (equation 2) of the popular Adam optimizer Kingma and Ba (2014). These are then concatenated in a block-diagonal way according to equation 3 to complete the process. After optimizing the relation parameters $\mathbf{e}_{R}\in\mathbb{R}^{n\times n}$, they are held constant while the entity parameters $\mathbf{e}_{V}\in\mathbb{R}^{n\times m}$ are optimized using Stochastic Gradient Descent (SGD), specifically employing the Adagrad optimizer Duchi et al. (2011).
5 Experiment
We expect that our proposed OrthogonalE model, employing matrices for entities and block-diagonal orthogonal matrices with Riemannian optimization for relations, will outperform baseline models. Also, we anticipate OrthogonalE is a general and flexible KGE model that saves parameters while effectively balancing high-dimensional block-diagonal orthogonal matrices $\mathbf{X}_{i}$ and entity matrices $\mathbf{e}_{V}$. Our goal is to validate these through empirical testing.
5.1 Experiment Setup
Dataset Entities Relations Train Validation Test
WN18RR 40,943 11 86,835 3,034 3,134
FB15k-237 14,541 237 272,115 17,535 20,466
Table 1: Details of the two datasets.
WN18RR FB15K-237
Model MRR H@1 H@3 H@10 MRR H@1 H@3 H@10
TransE $\diamondsuit$ .226 - - .501 .294 - - .465
DistMult $\diamondsuit$ .430 .390 .440 .490 .241 .155 .263 .419
ComplEx $\diamondsuit$ .440 .410 .460 .510 .247 .158 .275 .428
ConvE $\diamondsuit$ .430 .400 .440 .520 .325 .237 .356 .501
RotatE $\diamondsuit$ .470 .422 .488 .565 .297 .480 .328 .205
QuatE $\diamondsuit$ .481 .436 .500 .564 .311 .221 .342 .495
Gram-Schmidt($\times2$) .475 .434 .489 .556 .317 .226 .344 .502
Gram-Schmidt($\times3$) .487 .445 .500 .568 .322 .232 .350 .504
OrthogonalE($\times2$) .490 .445 .503 .573 .330 .239 .368 .516
OrthogonalE($\times3$) .493 .450 .508 .580 .331 .240 .359 .513
Table 2: Link prediction accuracy results of two datasets, Bold indicates the best score, and underline represents the second-best score. For a fair comparison, we standardized $m$ at 1 for Gram-Schmidt and all OrthogonalE sizes. The entity dimension for WN18RR was set to approximately 500 (e.g. 501 for $\times3$ blocks to ensure experimental feasibility), and around 1000 for FB15K-237. $\diamondsuit$]: Results are taken from Zhang et al. (2019). The result of RotatESun et al. (2019) and QuatEZhang et al. (2019) are reported without self-adversarial negative sampling and type-constrains respectively for fair comparison.
Dataset.
We evaluate our proposed method on two KG datasets, including WN18RR Dettmers et al. (2018), FB15K-237 Toutanova and Chen (2015). The details of these datasets are shown in Table 1.
Evaluation metrics.
To predict the tail entity from a given head entity and relation, we rank the correct tail entity among all possible entities using two established ranking metrics. First is the Mean Reciprocal Rank (MRR), the average inverse ranking of correct entities, calculated as $\frac{1}{n}\sum_{i=1}^{n}\frac{1}{\text{ Rank }_{i}}$. Second is Hits@K for $K\in\{1,3,10\}$, the frequency of correct entities ranking within the top $K$ positions.
Baselines.
We compare our new model with state-of-the-art (SOTA) methods, namely TransE Bordes et al. (2013), DistMult Yang et al. (2014), ComplEx Trouillon et al. (2016), ConvE Dettmers et al. (2018), RotatE Sun et al. (2019), and QuatE Zhang et al. (2019). In addition to these five and our OrthogonalE, we introduce comparative models utilizing the Gram-Schmidt process for generating orthogonal matrices and SGD for joint relation-entity training. OrthogonalE further differentiates by employing orthogonal matrices of varying sizes to discuss performance nuances.
Implementation
The key hyperparameters in our implementation include the learning rate for RiemannianAdam Kochurov et al. (2020) and Adagrad Duchi et al. (2011), negative sample size, and batch size. To determine the optimal hyperparameters, we performed a grid search using the validation data.
Figure 4: MRR accuracy comparison of OrthogonalE models with different small orthogonal matrices across varying entity dimensions (n1n*1, we set $m=1$) on WN18RR.
Figure 5: MRR accuracy comparison of OrthogonalE models with different small orthogonal matrices across varying entity dimensions (n1n*1, we set $m=1$) on FB15K-237.
Figure 6: MRR accuracy comparison of OrthogonalE($\times2$) and Gram-Schmidt($\times2$) models across varying entity dimensions ($m$) with fixed relation matrix ($\times40$) on WN18RR.
Figure 7: MRR accuracy comparison of RotatE and OrthogonalE($\times2$) models across varying entity dimensions (nmn*m) on WN18RR.
Figure 8: MRR accuracy comparison of QuatE and OrthogonalE($\times3$) models across varying entity dimensions (nmn*m) on WN18RR.
Figure 9: MRR accuracy comparison of RotatE and OrthogonalE($\times2$) models across varying entity dimensions (nmn*m) on FB15K-237.
5.2 Results
We first analyzed the overall accuracy for all baseline models and OrthogonalE, then separately examined the impacts of Block-diagonal Orthogonal matrices, Riemannian Optimization for relations, and Entity matrices on the model from various experimental results. These results provided evidence supporting OrthogonalE’s generality and flexibility.
Overall Accuracy
Table 2 presents link prediction accuracies for the WN18RR and FB15K-237 datasets. The OrthogonalE model surpasses all state-of-the-art models, highlighting its superior representational ability by employing matrices for entities and block-diagonal orthogonal matrices with Riemannian optimization for relations. Moreover, the OrthogonalE model with $\times2$ and $\times3$ configurations yields significantly better performance than corresponding sizes of the Gram-Schmidt method, and notably exceeds RotatE and QuatE, respectively, showcasing the enhanced efficacy of the KGE model.
Block-diagonal Orthogonal matrices
Figure 4 and 5 show MRR accuracy comparison of OrthogonalE models with different block-diagonal orthogonal matrices across varying entity dimensions (n1n*1, we set $m=1$) on WN18RR and FB15K-237, repectively.
An initial analysis of the datasets reveals that WN18RR contains 40,943 entities but only 11 relations, averaging 3,722 entities per relation. Conversely, FB15K-237 has 14,541 entities and 237 relations, averaging only 61 entities per relation. This suggests that WN18RR, in comparison to FB15K-237, requires a more complex and effective representation capability.
Our experimental findings align with the dataset analysis. For WN18RR, as shown in Figure 4, there is little difference in performance between $\times3$, $\times4$, $\times5$, and $\times10$ blocks, but these configurations significantly outperform the $\times2$ blocks. In contrast, for FB15K-237 depicted in Figure 5, the results across block-diagonal sizes, including $\times2$, $\times3$, $\times4$, and $\times10$ are consistent. This suggests that $\times2$ blocks suffice for representing FB15K-237’s relations, while WN18RR requires higher-dimensional, more complex blocks, such as $\times3$ or $\times4$, for adequate representation. These results illustrate that the OrthogonalE model is general, which can adapt to datasets of various complexities by adjusting the dimension $d$ of the block-diagonal matrices.
Riemannian Optimization for relations
Figure 6 compares MRR accuracies of OrthogonalE ($\times2$) and Gram-Schmidt ($\times2$) across entity dimensions ($m$) with a constant relation matrix ($\times40$) on WN18RR, assessing the efficacy of orthogonal optimization beyond the Gram-Schmidt method for block-diagonal orthogonal matrices. The result demonstrates that OrthogonalE’s Riemannian optimization significantly exceeds Gram-Schmidt, underscoring its necessity.
Entity matrix
Figures 7 and 8 compare MRR accuracies of RotatE with OrthogonalE ($\times2$) and QuatE with OrthogonalE ($\times3$) over different entity dimensions nmn*m on WN18RR, then Figure 9 shows comparison of RotatE and OrthogonalE ($\times2$) in FB15-237. In OrthogonalE, we maintained a constant entity dimension (nmn*m) while varying $m$ to assess the impact of entity shape. Figures 7 and 9 show that models with mm=1, 2, or 3(WN18RR) perform similarly and outperform $m=10$; notably, even when the number of relation parameters is just 1/10th of RotatE’s, the performance is significantly better at all dimensions. Figure 8 mirrors these findings, with mm=1 and 2 outperform mm=10, and mm=10 again surpassing QuatE, which has 7.5 times more parameters. These results demonstrate OrthogonalE’s efficacy in saving relation parameters while outperforming RotatE and QuatE, highlighting our model’s flexibility in controlling entity dimension through variable $m$ without unnecessarily increasing relation size.
Furthermore, Figure 6 result (with $m=7$ yielding MRR=0.483) suggests that a relation matrix of $\times40$ (20 parameters), compared to a dimension of 500 (250 parameters), can achieve comparably high performance, thus demonstrating that entity matrix method significantly reduces the need for excessive relation parameters to yield effective outcomes.
6 Conclusion
In this study, we propose the OrthogonalE model to acquire a flexible and general KGE model with employing matrices for entities and block-diagonal orthogonal matrices with Riemannian optimization for relations. The experimental results demonstrate that our new KGE model, OrthogonalE, is both general and flexible, significantly outperforming state-of-the-art KGE models while substantially reducing the number of relation parameters.
Acknowledgements